# Solvability of fractional semilinear heat equations with distributional forcing terms in Lorentz spaces

### v1

@@ 1.2 Main results
Let $\gamma>1$, $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$.
@@
    There exist positive constants $\delta$, $M$ and $T$ such that for every $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta}_{\left(p,\infty\right),1}\right\|<\delta$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ with a bound
    $\sup_{t\in\left]0,T\right[}\left\|u(x,t)\mid L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right\|\leqslant M.$ (1.3)
@@
    $\mu\in B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}(\mathbb{R}^{N})\quad\text{and}\quad\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}(\mathbb{R}^{N})\right\|<\delta$
    for some $\varepsilon\in\left]0,\theta\right[$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.3).
@@
    Assume $\theta<s<0$. There exists $M>0$ such that for every $\mu\in B^{s}_{(p,\infty),\infty}(\mathbb{R}^{N})$, problem (1.1) possesses a solution $u(x,t)$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.3). ∎
Each of these three sufficient conditions on $\mu$ has its own advantages. Note that $p=\frac{N(\gamma-1)}{\theta}$ is feasible only when the Serrin supercritical condition $\gamma>\frac{N}{N-\theta}$ holds, because we imposed $p>\gamma$.
@@
From Proposition 2.9 and Proposition 2.10, we have
$L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}\subset B^{0}_{\left(\frac{N(\gamma-1)}{\gamma\theta},\infty\right),\infty}\subset B^{\varepsilon-\theta}_{\left(\frac{N(\gamma-1)}{\theta+(\gamma-1)\varepsilon},\infty\right),\infty}$
for some $\varepsilon>0$. Thus in the Serrin supercritical case $\gamma>\frac{N}{N-\theta}$, by applying Theorem 1.5(2) with $p=\frac{N(\gamma-1)}{\theta}$, we find out that problem (1.1) possesses a solution for $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\mathrm{ul}}\right\|<\delta$. For example, we can take homogeneous functions $c|x|^{-\frac{\gamma\theta}{\gamma-1}}$ or their infinite sum
@@
as forcing terms for sufficiently small $c>0$. This is compatible with the result of [11]. Moreover, the embeddings above are non-trivial, our sufficient condition is slightly broader than their results.
@@
We know that all $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta}}(\mathbb{R}^{N})$ are acceptable as forcing term. Recall that problem (1.1) has the following structure: If a funciton $u(x,t)$ is a solution to problem (1.1) with a forcing term $\mu(x)$, then a function $u_{\lambda}(x,t):=\lambda^{\frac{\theta}{\gamma-1}}u(\lambda x,\lambda^{\theta}t)$ is a solution to problem (1.1) with $\mu_{\lambda}(x):=\lambda^{\frac{\gamma\theta}{\gamma-1}}\mu(\lambda x)$. For any $\mu\in L^{\frac{N(\gamma-1)}{\gamma\theta}}(\mathbb{R}^{N})$, we see that
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu_{\lambda}\mathrel{}\middle|\mathrel{}L^{\frac{Np}{N+p\varepsilon},\infty}_{\mathrm{ul}}\right\|$ $\leqslant\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta}_{\left(\frac{Np}{N+p\varepsilon},\infty\right),\infty}\right\|\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{0}_{\left(\frac{N(\gamma-1)}{\gamma\theta},\infty\right),\infty}\right\|$
$\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}L^{\frac{N(\gamma-1)}{\gamma\theta},\infty}_{\text{ul}}\right\|$ $\leqslant C\sup_{z\in\mathbb{R}^{N}}\left(\int_{|x-z|\leqslant 1}\lambda^{N}|\mu(\lambda x)|^{\frac{N(\gamma-1)}{\gamma\theta}}\,dx\right)^{\frac{\gamma\theta}{N(\gamma-1)}}$
@@
holds for all $\lambda>0$. By absolute continuity of $\left|\mu\right|^{\frac{N(\gamma-1)}{\gamma\theta}}$, the last term tends to zero if we send $\lambda\to 0$. Thus we can apply Remark 1.6 to $\mu_{\lambda}$, $0<\lambda\ll 1$.
@@
In the Serrin subcritical case $\gamma<\frac{N}{N-\theta}$, we can pick $p_{0}$ as $\gamma<p_{0}<\frac{N}{N-\theta}$. Then by Sobolev embedding (Proposition 2.10), we have
$L^{1}\subset B^{0}_{1,\infty}\subset B^{N/p_{0}-N}_{p_{0},\infty}\subset B^{N/p_{0}-N}_{(p_{0},\infty),\infty}$
and the space appearing in the last satisfies the assumptions of Theorem 1.5 (3). For instance, we know that the Dirac delta measure $\delta(x)$ belongs to $B^{0}_{1,\infty}\subset B^{N/p_{0}-N}_{p_{0},\infty}\subset B^{N/p_{0}-N}_{(p_{0},\infty),\infty}$. Furthermore, if $\theta\in\left]1,2\right]$ and $\gamma>1$ satisfies $\gamma<\frac{N}{N+1-\theta}$, we can take $\partial_{x_{i}}\delta(x)$ as a forcing term of problem (1.1) since we can take $p_{1}\in\left]\gamma,\frac{N}{N+1-\theta}\right[$ and we have
$\partial_{x_{i}}\delta(x)\in B^{N/p_{1}-N-1}_{p_{1},\infty}\subset B^{N/p_{1}-N-1}_{(p_{1},\infty),\infty}\quad\text{for}\quad i\in\left\{1,...,N\right\}.$
This example reflects a benefit of employing Besov-type spaces. We also note that employing Besov–Morrey spaces (see Definition 1.2 in [30]) can broaden the class of acceptable forcing terms in this case ($\gamma<N/(N-\theta)$).
In this paper, we do not treat in detail the Serrin critical case $\gamma=\frac{N}{N-\theta}$. This case requires some special function spaces (see Theorem 1.3 in [11] for details).
The next theorem states a necessary condition on forcing terms $\mu$ for the existence of a solution $u=u(x,t)$ belonging to $L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$.
@@
Let $p>\gamma$ and $p\geqslant\frac{N(\gamma-1)}{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}(0,T;L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N}))$ satisfies the integral equation (1.2) for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. Then, the tempered distribution $\mu$ appering in (1.2) must be an element of $B^{-\theta}_{(p,\infty),\infty}(\mathbb{R}^{N})$. ∎
@@
We claim nothing about existence of solutions to problem (1.1) for $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying
[…]
By similarly carrying the arguments of the proof of Theorem 1.9 with the aid of Theorem 2 in [1] and of the Hardy–Littlewood–Sobolev inequality (see also [16]), we can obtain a necessary condition on the solvability of the following Cauchy problem
@@
where $\gamma>1$, $\theta\in\left]0,2\right]$.
@@
for almost all $(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$ for some $T>0$. Then, the tempered distribution $\mu(x)$ appering in (1.5) must be an element of inhomogeneous Besov space $B^{-\theta/q}_{p,q}(\mathbb{R}^{N})=B^{N/p-\theta/(\gamma-1)}_{p,\theta p(\gamma-1)/(\theta p-N\gamma+N)}(\mathbb{R}^{N})$. ∎
The rest of this paper proceeds as follows. We divide Section 2 into five subsections. The first three are devoted to collecting basic facts, including real interpolation for Lorentz and Besov spaces. We recall decay estimates of operators $S(t)$ in the next subsection. At the end of Section 2, we make an estimate of the Besov–Lorentz norm of the forcing term. When $\theta<2$, it needs additional care and we postpone this task to Appendix A. In Section 3, we prove Theorem 1.5 by applying the contraction mapping theorem and Theorem 1.9 by applying Proposition 2.16.
@@ 2 Preliminaries
In this section, we gather some properties of Lorentz spaces $L^{p,q}$, $L^{p,q}_{\mathrm{ul}}$, of Besov–Lorentz spaces $B^{s}_{(p,q),r}$ and of operators $S(t)$ on these spaces. We also show that Besov–Lorentz norms $\left\|f\mathrel{}\middle|\mathrel{}B^{-\theta}_{(p,q),\infty}\right\|$ can be estimated by a space-time norm of the Duhamel integral of $f$ (Proposition 2.16).
2.1 Basic properties of Lorentz spaces
We begin with elemental properties of Lorentz spaces.
@@ 2.1 Basic properties of Lorentz spaces
Fix $\mathcal{L}\in\left\{L,L_{\mathrm{ul}}\right\}$ and $p\in\left]0,\infty\right[$. The following claims hold.
@@
    Assume that $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p^{-1}=p_{0}^{-1}+p_{1}^{-1}$. Then, there exists $C>0$ depending on $p_{0}$, $p_{1}$ such that the inequality
    $\left\|fg\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|\leqslant C\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{p_{0},\infty}\right\|\left\|g\mathrel{}\middle|\mathrel{}\mathcal{L}^{p_{1},\infty}\right\|$
    holds for all $f\in\mathcal{L}^{p_{0},\infty}$ and $g\in\mathcal{L}^{p_{1},\infty}$.
@@
    Fix $r\in\left]0,\infty\right[$. Then, the relation
    $\left\||f|^{r}\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|=\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{pr,\infty}\right\|^{r}$
    holds for all $f\in\mathcal{L}^{pr,\infty}$.
@@
    Assume $|g(x)|\leqslant|f(x)|$ for a.e.-$x$. Then, we have
    $\left\|g\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}\mathcal{L}^{p,\infty}\right\|.$
@@
    $\left\|f_{\lambda}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|=\lambda^{\frac{-N}{p}}\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}\right\|$
    , where $f_{\lambda}(x):=f(\lambda x)$ and $q\in\left]0,\infty\right[$.
@@
(1) for $\mathcal{L}=L$: See Exercise 1.1.15 in [7]. Others are easily confirmed. ∎
We recall the Young inequality for Lorentz spaces.
@@
Let $p\in\left]1,\infty\right[$, $q\in\left[1,\infty\right]$. There exists a positive number $C$ depending on $p$ such that the inequality
$\left\|g\ast f\mid L^{p,q}(\mathbb{R}^{N})\right\|\leqslant C\left\|g\mid L^{1}(\mathbb{R}^{N})\right\|\left\|f\mid L^{p,q}(\mathbb{R}^{N})\right\|$
holds for all $g\in L^{1}(\mathbb{R}^{N})$ and $f\in L^{p,\infty}(\mathbb{R}^{N})$.
Proof.
See Theorem 2.6. in [23] (For $q=\infty$, it is enough to see Theorem 1.2.13 in [7] (pp. 23)). ∎
We prepare the Young inequality for uniformly local Lorentz spaces.
Proposition 2.3.
Let $p\in\left]1,\infty\right[$. There exists a positive constant $C$ depending on $p$ such that the inequality
$\left\|g\ast f\mid L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})\right\|\leqslant C\left\|g\mid L^{1}(\mathbb{R}^{N})\right\|\left\|f\mid L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})\right\|$
holds for all $g\in L^{1}(\mathbb{R}^{N})$ and $f\in L^{p,\infty}_{\text{ul}}(\mathbb{R}^{N})$.
It is worth mentioning that Propostion 2.3 does not directly follow from Proposition 2.2. We shall modify the proof of Proposition 2.2 (see pp. 24–25 in [7]) and provide the full proof here for completeness.
@@
Fix $z\in\mathbb{R}^{N}$ and $\lambda>0$. Take any $f\in L^{p,\infty}_{\mathrm{ul}}(\mathbb{R}^{N})$ and $g\in L^{1}(\mathbb{R}^{N})$. Set $M:=\frac{\lambda}{2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|}$. We decompose $f$ as
@@
Because $\left|f_{0}\ast g(x)\right|\leqslant\lambda/2$ holds for all $x\in\mathbb{R}^{N}$, we have
$\alpha_{(f\ast g)\chi_{B(z,1)}}(\lambda)\leqslant\alpha_{(f_{0}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)+\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right)=\alpha_{(f_{1}\ast g)\chi_{B(z,1)}}\left(\frac{\lambda}{2}\right).$

### v2

@@ 1.2 Main results
Let $\gamma>1$, $\gamma<q\leqslant p$ and $p\geqslant{N(\gamma-1)}/{\theta}$.
@@
    There exist positive constants $\delta$, $M$ and $T$ such that for every $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}B^{-\theta,1}_{p,q,\infty}\right\|<\delta$, problem (P) possesses a solution $u$ on $\left]0,T\right[$ with a bound
    $\sup_{t\in\left]0,T\right[}\left\|u(\cdot,t)\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}(\mathbb{R}^{N})\right\|\leqslant M.$ (1.2)
@@
    $\mu\in B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}(\mathbb{R}^{N})\quad\text{and}\quad\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu\mathrel{}\middle|\mathrel{}M^{\frac{Np}{N+p\varepsilon}}_{\frac{Nq}{N+p\varepsilon},\infty}(\mathbb{R}^{N})\right\|<\delta$
    for some $\varepsilon\in\left]0,\min\left\{\theta,N(q-1)/p\right\}\right[$, problem (P) possesses a solution $u$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.2).
@@
    Assume $\theta<s<0$. There exists $M>0$ such that for every $\mu\in B^{s,\infty}_{p,q,\infty}(\mathbb{R}^{N})$, problem (P) possesses a solution $u$ on $\left]0,T\right[$ for some $T>0$ with a bound (1.2). ∎
Each of these three sufficient conditions on $\mu$ has its own advantages. Note that $p={N(\gamma-1)}/{\theta}$ is feasible only when the Serrin supercritical condition $\gamma>\gamma_{\ast}$ holds, because we imposed $p\geqslant q>\gamma$.
@@
Let us consider the Serrin supercritical case $\gamma>\gamma_{\ast}={N}/{(N-\theta)}$ with $\theta<N$. From Proposition 2.11 and Proposition 2.12, we have
$M^{\frac{N(\gamma-1)}{\gamma\theta}}_{{q}/{\gamma}}\subset M^{\frac{N(\gamma-1)}{\gamma\theta}}_{{q}/{\gamma},\infty}\subset B^{0,\infty}_{\frac{N(\gamma-1)}{\gamma\theta},\frac{q}{\gamma},\infty}\subset B^{\varepsilon-\theta,\infty}_{\frac{N(\gamma-1)}{\theta+(\gamma-1)\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}$
for some $\varepsilon>0$. Thus, by applying Theorem 1.5 (2) with $p={N(\gamma-1)}/{\theta}$ and any $q\in\left]\gamma,p\right[$, we find out that problem (P) possesses a solution for $\mu\in M^{{N(\gamma-1)}/{\gamma\theta},\infty}_{q/\gamma}$ satisfying $\left\|\mu\mathrel{}\middle|\mathrel{}M^{{N(\gamma-1)}/{\gamma\theta},\infty}_{q/\gamma}\right\|<\delta$. For example, we can take homogeneous functions $c|x|^{-\frac{\gamma\theta}{\gamma-1}}$ or their infinite sum
@@
as inhomogeneous terms for sufficiently small $c>0$. This is compatible with the results of [12, 14]. See Remark 1.8 and Proposition B.1 for further examples; therein lies the advantage of employing Morrey-type spaces.
@@
Assume $\gamma>\gamma_{\ast}$. We know that all $\mu\in L^{{N(\gamma-1)}/{\gamma\theta}}(\mathbb{R}^{N})$ are acceptable as inhomogeneous terms. Recall that problem (P) has the following structure: If a function $u(x,t)$ is a solution to problem (P) with a inhomogeneous term $\mu(x)$, then a function $u_{\lambda}(x,t):=\lambda^{\frac{\theta}{\gamma-1}}u(\lambda x,\lambda^{\theta}t)$ is a solution to problem (P) with $\mu_{\lambda}(x):=\lambda^{\frac{\gamma\theta}{\gamma-1}}\mu(\lambda x)$. Let $p=N(\gamma-1)/\theta$. For any $\mu\in L^{{N(\gamma-1)}/{\gamma\theta}}(\mathbb{R}^{N})$, we see that
$\limsup_{j\to\infty}2^{(\varepsilon-\theta)j}\left\|\Delta_{j}\mu_{\lambda}\mathrel{}\middle|\mathrel{}M^{\frac{Np}{N+p\varepsilon}}_{\frac{Nq}{N+p\varepsilon},\infty}\right\|$ $\leqslant\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right\|\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}B^{0,\infty}_{\frac{N(\gamma-1)}{\gamma\theta},\frac{q}{\gamma},\infty}\right\|$
$\leqslant C\left\|\mu_{\lambda}\mathrel{}\middle|\mathrel{}M^{\frac{N(\gamma-1)}{\gamma\theta}}_{\frac{N(\gamma-1)}{\gamma\theta},\infty}\right\|$ $\leqslant C\sup_{z\in\mathbb{R}^{N}}\left(\int_{|x-z|\leqslant 1}\lambda^{N}|\mu(\lambda x)|^{\frac{N(\gamma-1)}{\gamma\theta}}\,dx\right)^{\frac{\gamma\theta}{N(\gamma-1)}}$
@@
holds for all $\lambda>0$. Then, by absolute continuity of $\left|\mu\right|^{\frac{N(\gamma-1)}{\gamma\theta}}$, the last term tends to zero if we send $\lambda\to 0$. Thus we can apply Remark 1.6 to $\mu_{\lambda}$, $0<\lambda\ll 1$.
@@
Assume $\theta>1$ and $\gamma>\gamma_{\ast}$. Set $\delta:=\frac{1+(\theta-1)\gamma}{N(\gamma-1)}$ and take any $q\in\left]\gamma,{N(\gamma-1)}/{\theta}\right[$. These conditions give $1<\frac{\theta q}{1+(\theta-1)\gamma}<1/\delta$. Then [20, Proposition 1.3], [20, Lemma 1.4] and Proposition 2.11 imply
$\mu(x):=\prod_{j=1}^{N}\left(x_{j}\right)_{+}^{-\delta}\in M^{1/\delta}_{\theta q/(1+(\theta-1)\gamma)}\subset B^{0,\infty}_{\frac{1}{\delta},\frac{\theta q}{1+(\theta-1)\gamma},\infty}$
and thus we see that
$\frac{1}{-\delta}\frac{\partial\mu}{\partial x_{1}}=\mathrm{f.p.}\,(x_{1})_{+}^{-\delta-1}\prod_{j=2}^{N}(x_{j})_{+}^{-\delta}\in B^{-1,\infty}_{\frac{1}{\delta},\frac{\theta q}{1+(\theta-1)\gamma},\infty}$
holds. Since the space $B^{-1,\infty}_{{1}/{\delta},{\theta q}/{\left(1+(\theta-1)\gamma\right)},\infty}$ satisfies the assumptions of Theorem 1.5 (2) with $\varepsilon=\theta-1$ and $p=N(\gamma-1)/\theta$, it turns out that the distribution $c(\partial\mu/\partial x_{1})$ can be taken as the inhomogeneous term if $0<c\ll 1$. We also note that $\mu$ does not belong to Lorentz spaces if $N\geqslant 2$.
Remark 1.9.
In the Serrin subcritical case $\gamma<\gamma_{\ast}$, we can pick $p_{0}$ as $\gamma<p_{0}<\gamma_{\ast}$. Then by Proposition 2.12, we have
$L^{1}\subset B^{0,\infty}_{1}\subset B^{N/p_{0}-N,\infty}_{p_{0}}\subset B^{N/p_{0}-N,\infty}_{p_{0},p_{0},\infty}$
and the space appearing in the last satisfies the assumptions of Theorem 1.5 (3). For instance, we know that the Dirac delta measure $\delta(x)$ belongs to $B^{0,\infty}_{1}\subset B^{N/p_{0}-N,\infty}_{p_{0}}\subset B^{N/p_{0}-N,\infty}_{p_{0},p_{0},\infty}$. Furthermore, if $\theta>m$ with $m\in\mathbb{Z}_{\geqslant 1}$ and if $\gamma>1$ satisfies $\gamma<{N}/{(N+m-\theta)}$, we can take $\partial^{\alpha}\delta(x)$ as an inhomogeneous term of problem (P) since we can take $p_{1}\in\left]\gamma,{N}/{(N+m-\theta)}\right[$ and we have
$\partial^{\alpha}\delta(x)\in B^{N/p_{1}-N-m,\infty}_{p_{1}}\subset B^{N/p_{1}-N-m,\infty}_{p_{1},p_{1},\infty}\quad\text{for }i\in\left\{1,...,N\right\}\,\,\text{and for }|\alpha|\leqslant m.$
This example reflects a benefit of employing Besov-type spaces.
In this paper, we do not treat in detail the Serrin critical case $\gamma=\gamma_{\ast}$. This case requires some special function spaces (see Theorem 1.3 in [14] for details).
The next theorem states a necessary condition on inhomogeneous terms $\mu$ for the existence of a solution $u=u(x,t)$ belonging to $L^{\infty}(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N}))$.
@@
Let $T>0$, $p\geqslant q>\gamma$ and $p\geqslant{N(\gamma-1)}/{\theta}$. Assume that a function $u=u(x,t)\in L^{\infty}\left(0,T;M^{p}_{q,\infty}(\mathbb{R}^{N})\right)$ satisfies the integral equation (1.1) for a.e.-$(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$. Then the tempered distribution $\mu$ appearing in (1.1) must be an element of $B^{-\theta,\infty}_{p,q,\infty}(\mathbb{R}^{N})$. ∎
@@
We claim nothing about existence of solutions to problem (P) for $\mu\in\mathcal{S}^{\prime}(\mathbb{R}^{N})$ satisfying
$\mu\in B^{-\theta,\infty}_{p,q,\infty}\setminus\left\{B^{-\theta,1}_{p,q,\infty}\cup\left(\bigcup_{0<\varepsilon<\theta}B^{\varepsilon-\theta,\infty}_{\frac{Np}{N+p\varepsilon},\frac{Nq}{N+p\varepsilon},\infty}\right)\right\}.$
By similarly carrying the arguments of the proof of Theorem 1.10 with the aid of Theorem 2 in [1] and of the Hardy–Littlewood–Sobolev inequality (see also [19]), we can obtain a necessary condition on the initial data for the local-in-time solvability of the following Cauchy problem
@@
where $\gamma>1$, $\theta>0$.
@@
for a.e.-$(x,t)\in\mathbb{R}^{N}\times\left]0,T\right[$ for some $T>0$. Then the tempered distribution $\mu(x)$ appearing in (1.4) must be an element of inhomogeneous Besov space $B^{-\theta/q,q}_{p}(\mathbb{R}^{N})=B^{N/p-\theta/(\gamma-1),q}_{Nq(\gamma-1)/\theta(q-\gamma+1)}(\mathbb{R}^{N})$. ∎
The rest of this paper proceeds as follows. We divide Section 2 into five subsections. The first three are devoted to establishing basic properties, including real interpolation for Lorentz–Morrey spaces and Besov spaces. We confirm decay estimates of operators $S(t)$ in the next subsection. At the end of Section 2, we make an estimate of the BLM norm of the inhomogeneous term. When $\theta\neq 2$, it needs additional care and we postpone this task to Appendix A. In Section 3, we prove Theorem 1.5 by applying the contraction mapping theorem and Theorem 1.10 by applying Proposition 2.18. We make some comments on Morrey spaces in Appendix B.
@@ 2 Preliminaries
In this section, we gather some preliminary properties of Lorentz spaces $L^{p,q}$, of Lorentz–Morrey spaces $M^{p}_{q,r}$ and of BLM spaces $B^{s,r}_{p,q,\kappa}$. We also confirm decay estimates of operators $S(t)$ on these spaces. In the last subsection, we show that BLM norms $\left\|f\mathrel{}\middle|\mathrel{}B^{-\theta,\infty}_{p,q,\kappa}\right\|$ can be estimated by a space-time norm of the Duhamel integral of $f$.
2.1 Basic properties of Lorentz–Morrey spaces
We begin with recalling basic properties of Lorentz spaces.
@@ 2.1 Basic properties of Lorentz–Morrey spaces
Fix $p\in\left]0,\infty\right[$. The following claims hold.
@@
    Assume that $p_{0}$ and $p_{1}\in\left]0,\infty\right[$ satisfy $p^{-1}=p_{0}^{-1}+p_{1}^{-1}$. Then there exists $C>0$ depending on $p_{0}$, $p_{1}$ such that the inequality
    $\left\|fg\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|\leqslant C\left\|f\mathrel{}\middle|\mathrel{}L^{p_{0},\infty}\right\|\left\|g\mathrel{}\middle|\mathrel{}L^{p_{1},\infty}\right\|$
    holds for all $f\in L^{p_{0},\infty}$ and $g\in L^{p_{1},\infty}$.
@@
    Fix $r\in\left]0,\infty\right[$. Then the identity
    $\left\||f|^{r}\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|=\left\|f\mathrel{}\middle|\mathrel{}L^{pr,\infty}\right\|^{r}$
    holds for all $f\in L^{pr,\infty}$.
@@
    Assume $|g(x)|\leqslant|f(x)|$ for a.e.-$x$. Then we have
    $\left\|g\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}L^{p,\infty}\right\|.$
@@
    $\left\|f_{\lambda}\mathrel{}\middle|\mathrel{}L^{p,q}\right\|=\lambda^{\frac{-N}{p}}\left\|f\mathrel{}\middle|\mathrel{}L^{p,q}\right\|,$
    where $f_{\lambda}(x):=f(\lambda x)$ and $q\in\left]0,\infty\right]$.
@@
(1) See Exercise 1.1.15 in [9]. Others are easily confirmed. ∎
Lemma 2.1 leads the corresponding properties of Lorentz–Morrey spaces.
Lemma 2.2.
Fix $0<q\leqslant p<\infty$. The following claims hold.
  1. (1)
    Assume that $p_{0}$, $p_{1}$, $q_{0}$ and $q_{1}\in\left]0,\infty\right[$ satisfy $p_{0}>q_{0}$, $p_{1}>q_{1}$ and $\left(p^{-1},q^{-1}\right)=\left(p_{0}^{-1}+p_{1}^{-1},q_{0}^{-1}+q_{1}^{-1}\right)$. Then there exists $C>0$ depending on $p_{0}$, $p_{1}$ $q_{0}$ and $q_{1}$ such that the inequality
    $\left\|fg\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant C\left\|f\mathrel{}\middle|\mathrel{}M^{p_{0}}_{q_{0},\infty}\right\|\left\|g\mathrel{}\middle|\mathrel{}M^{p_{1}}_{q_{1},\infty}\right\|$
    holds for all $f\in M^{p_{0}}_{q_{0},\infty}$ and $g\in M^{p_{1}}_{q_{1},\infty}$.
  2. (2)
    Fix $r\in\left]0,\infty\right[$. Then the identity
    $\left\||f|^{r}\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|=\left\|f\mathrel{}\middle|\mathrel{}M^{pr}_{qr,\infty}\right\|^{r}$
    holds for all $f\in M^{pr}_{qr,\infty}$.
  3. (3)
    Assume $|g(x)|\leqslant|f(x)|$ for a.e.-$x$. Then we have
    $\left\|g\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|\leqslant\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}\right\|.$
We prepare the Young inequality for Lorentz–Morrey spaces.
@@
Let $1<q\leqslant p<\infty.$ There exists a positive constant $C$ depending on $q$ such that the inequality
$\left\|g\ast f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}(\mathbb{R}^{N})\right\|\leqslant C\left\|g\mathrel{}\middle|\mathrel{}L^{1}(\mathbb{R}^{N})\right\|\left\|f\mathrel{}\middle|\mathrel{}M^{p}_{q,\infty}(\mathbb{R}^{N})\right\|$
holds for all $g\in L^{1}(\mathbb{R}^{N})$ and $f\in M^{p}_{q,\infty}(\mathbb{R}^{N})$.
Proposition 2.3 does not seem to directly follow from its counterpart of Lorentz spaces. We provide a proof here for completeness; the following argument is a modification of that on pp. 24–25 in [9].
@@
Fix any $R\in\left]0,1\right]$, $z\in\mathbb{R}^{N}$ and $\lambda>0$. Take arbitrary $f\in M^{p}_{q,\infty}(\mathbb{R}^{N})$ and $g\in L^{1}(\mathbb{R}^{N})$. Set $M:=\lambda/2\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|$. We decompose $f$ as
@@
Because $\left|f_{0}\ast g(x)\right|\leqslant M\left\|g\mathrel{}\middle|\mathrel{}L^{1}\right\|=\lambda/2$ holds for all $x\in\mathbb{R}^{N}$, we have
$\alpha_{(f\ast g)\chi_{B(z,R)}}(\lambda)\leqslant\alpha_{(f_{0}\ast g)\chi_{B(z,R)}}\left(\frac{\lambda}{2}\right)+\alpha_{(f_{1}\ast g)\chi_{B(z,R)}}\left(\frac{\lambda}{2}\right)=\alpha_{(f_{1}\ast g)\chi_{B(z,R)}}\left(\frac{\lambda}{2}\right).$